find the potential gradient between two parallel plate conductor of separation 0.03 when potential difference of 540 supplied to them hence determine the force in an electron which club between the plates

1
Expert's answer
2022-04-14T11:18:33-0400

The potential gradient between two parallel plate conductor

"E=\frac{V}{d}=\frac{540\:\rm V}{0.03\:\rm m}=18000\:\rm V/m"

The force acting on an electron

"F=qE\\
=\rm 1.6*10^{-19}\: C\times 18000\:\rm V/m=2.88*10^{-15}\: N"